# Fair Minimum Labeling: Efficient Temporal Network Activations for Reachability and Equity

### v1

@@ 5.2 A bicriterial approximation
The running time of $\mathcal{O}(n^{5})$ of the exact algorithm for solving the FML problem in trees limits its scalability. The bottleneck is the linear number of labels per node, and the pairwise merging costing $\mathcal{O}(n^{4})$ per node. So the idea is to reduce the number of labels per node in a principled way. A classic method to reduce dynamic programming overhead is label sparsification by grid rounding. The key is to group labels with similar $(b,r)$ values into the same bucket $B(\cdot,\cdot)$ and keep only the best, i.e., lowest cost $c$, label in that bucket.
@@ 5.3 Projecting the solution to the graph G
Let $\lambda_{T}$ be the temporal labeling obtained by one of these algorithms on the tree $T$. The tree labeling $\lambda_{T}$ must be translated into a valid temporal labeling $\lambda_{G}$ for the original graph $G$. This projection ensures that the reachability achieved in the tree is preserved in the graph. For each temporal tree edge activation $(\{u,v\},t)\in\lambda_{T}$ we first identify a shortest path $P_{uv}$ between $u$ and $v$ in the original graph $G$. Let this path consist of the sequence of $p=d_{G}(u,v)$ edges $e_{1},e_{2},\dots,e_{p}$. To mimic the activation of the tree edge $\{u,v\}$ at time $t$, we activate the path $P_{uv}$ in $G$ using a sequence of timestamps. The final graph labeling $\lambda_{G}$ is the union of all temporal edges generated by projecting all $(\{u,v\},t)\in\lambda_{T}$.
@@ 5.4 Generalizations
The algorithms and the specific complexity analysis presented in the previous sections are for the two-group (plus no-color) case. However, the framework is generalizable. For $k$ distinct groups, the dynamic programming label becomes $(g_{1},g_{2},\ldots,g_{k},c)$, where $g_{i}$ is the count of covered nodes from group $i$. The state space for the counts becomes $O(n^{k})$. The pairwise merge operation takes $O((n^{k})^{2})=O(n^{2k})$ time. With $n$ nodes, the total runtime for the exact DP is $O(n^{2k+1})$.
Our framework can also be extended to support non-uniform activation costs $w(e)$, while preserving the $O(\log n)$ approximation guarantee, by defining the metric for the FRT embedding using these costs. First, we compute a shortest-path metric $d(u,v)$ where the distance between nodes is defined as the minimum cost of a path between them, using the activation costs $w(e)$ as the edge weights. We then use FRT to embed this weighted metric into a tree $T$. Our DP algorithm then runs on the tree $T$, and because the tree distances being optimized are already low-distortion approximations of the minimum activation costs, the $O(\log n)$ guarantee on the final, total weighted cost holds directly.
Finally, in dynamic environments (e.g., node failures or departures), partial recomputation of our labeling is feasible. The DP can be rerun locally on affected subtrees, and minor changes in the graph structure often preserve the validity of the tree embedding. While full dynamic adaptation is beyond the scope of our current work, the methods lend themselves to such extensions.
@@ 6 Experiments
  • Closest greedily activates unweighted shortest paths from the sink to the closest uncovered colored node (regardless of group) until $\alpha$-coverage is achieved for both groups.
@@ 6.1.1 Network topology and data sources
We simulate a networked learning scenario over a random geometric graph $G=(V,E)$ with $|V|=1024$ nodes and connection radius $r=0.2$. The designated terminal node $t\in V$ collects data from a subset of data sources $B\cup R\subseteq V\setminus\{s\}$. Each of the data sources belongs to one of two demographic groups $B$ or $R$, assigned based on spatial proximity to the sink (near vs. far).
@@
The algorithms activate edges $A\subseteq E$ to connect data sources to the terminal. For the baselines, the activation cost is defined as the sum of hop-distances from all covered data sources to the sink. The collected data is used to train a shared MLP classifier. We use a fully connected MLP using an input layer of size $d=2$, followed by two hidden layers with 32 neurons each and ReLU activations, and a final output layer with a single neuron and sigmoid activation for binary classification. We train the model using the Adam optimizer [22] with learning rate $0.01$, batch size $64$, and weight decay $10^{-2}$ for $\ell_{2}$ regularization. The loss is binary cross-entropy. Training runs for 100 epochs with early stopping based on validation loss. The model is trained using the combined dataset collected from data sources temporally connected to the sink. Accuracy is reported separately on held-out test sets for each group. We also report the coverage of nodes in $B$ and $R$ that can temporally reach $t$. We report the mean values and standard deviations over ten independent runs.
@@ 6.3 Practical scalability
We sampled a connected 20,000-node induced subgraph (201,900 edges) from the Pokec dataset [32]. We randomly selected one node as the terminal and colored 10% of the remaining nodes (balanced 50/50) using the binary gender attribute (Blue = male, Red = female). We then ran FMLbiApprox with full coverage requirement ($\alpha=1.0$) and two precision settings ($\varepsilon=0.01$ and $\varepsilon=0.001$) in 10 independent runs. Table˜2 shows the results. Accuracy represents the ratio of achieved coverage to the required coverage. This real-world experiment confirms the practical scalability and effectiveness of our FMLbiApprox algorithm, which achieves high accuracy under both precision settings.
@@ 7 Limitations
Technical limitations. Our approximation algorithms currently focus on the case of a single terminal and two groups. While this setting already models common use cases, generalizing to arbitrary numbers of terminals and group types remains open. The exact dynamic programming algorithm has a running time complexity in $\mathcal{O}(n^{5})$, limiting scalability. The bicriteria approximation improves efficiency, reducing runtime to $\mathcal{O}(n^{2}+n\varepsilon^{-4}\log^{4}n)$, but trades off exact fairness for scalability: its guarantees degrade with the tree height, leading to a $(1+\varepsilon)^{(\log n)}$ factor in group coverage.
@@ 15     Keep only the label with minimum $c$ for each $(b,r)$ in $\mathrm{Labels}(v)$
Appendix A Omitted Proof
@@
The optimal weighted cost on the tree, $C_{T,opt}$, must satisfy $C_{T,opt}\leq C_{T}(\lambda_{G}^{*})$, as $\lambda_{G}^{*}$ provides one way to achieve the required reachability on $T$.
Our tree algorithm finds a solution $\lambda_{T}$ with weighted cost $C_{T}$ such that $C_{T}\leq C_{T,opt}$, either exactly, or via the approximation. Therefore, it follows $|\lambda_{G}|\leq C_{T}\leq C_{T,opt}\leq C_{T}(\lambda_{G}^{*})$.

### v2

@@ 5.2 A bicriterial approximation
The running time of $\mathcal{O}(n^{5})$ of the exact algorithm for solving the FML problem in trees limits its scalability. The bottleneck is the linear number of labels per node, and the pairwise merging costing $\mathcal{O}(n^{4})$ per node. So the idea is to reduce the number of labels per node in a principled way. A classic method to reduce dynamic programming overhead is label sparsification by grid rounding. The key is to group labels with similar $(b,r)$ values into a common bucket and keep only the best, i.e., lowest cost $c$, label in that bucket.
@@ 5.3 Projecting the solution to the graph G
Let $\lambda_{T}$ be the temporal labeling obtained by one of these algorithms on the tree $T$. The tree labeling $\lambda_{T}$ must be translated into a valid temporal labeling $\lambda_{G}$ for the original graph $G$. This projection ensures that the reachability achieved in the tree is preserved in the graph. For each temporal tree edge activation $(\{u,v\},\tau)\in\lambda_{T}$ we first identify a shortest path $P_{uv}$ between $u$ and $v$ in the original graph $G$. Let this path consist of the sequence of $p=d_{G}(u,v)$ edges $e_{1},e_{2},\dots,e_{p}$. To mimic the activation of the tree edge $\{u,v\}$ at time $\tau$, we activate the path $P_{uv}$ in $G$ using a sequence of timestamps. The final graph labeling $\lambda_{G}$ is the union of all temporal edges generated by projecting all $(\{u,v\},\tau)\in\lambda_{T}$.
@@ 5.4 Generalizations
The algorithms and the specific complexity analysis presented in the previous sections are for the two-group (plus no-color) case. However, the framework is generalizable. For $k$ distinct groups, the dynamic programming label becomes $(g_{1},g_{2},\ldots,g_{k},c)$, where $g_{i}$ is the count of covered nodes from group $i$. The state space for the counts becomes $\mathcal{O}(n^{k})$. The pairwise merge operation takes $\mathcal{O}((n^{k})^{2})=\mathcal{O}(n^{2k})$ time. With $n$ nodes, the total runtime for the exact DP is $\mathcal{O}(n^{2k+1})$.
Our framework can also be extended to support non-uniform activation costs $w(e)$, while preserving the $\mathcal{O}(\log n)$ approximation guarantee, by defining the metric for the FRT embedding using these costs. First, we compute a shortest-path metric $d(u,v)$ where the distance between nodes is defined as the minimum cost of a path between them, using the activation costs $w(e)$ as the edge weights. We then use FRT to embed this weighted metric into a tree $T$. Our DP algorithm then runs on the tree $T$, and because the tree distances being optimized are already low-distortion approximations of the minimum activation costs, the $\mathcal{O}(\log n)$ guarantee on the final, total weighted cost holds directly.
Finally, in dynamic environments (e.g., node arrivals, failures or departures), partial recomputation of our labeling is feasible. The DP can be rerun locally on affected subtrees, and minor changes in the graph structure often preserve the validity of the tree embedding. While full dynamic adaptation is beyond the scope of our current work, the methods lend themselves to such extensions.
@@ 6 Experiments
  • Closest greedily activates unweighted shortest paths from the terminal to the closest uncovered colored node (regardless of group) until $\alpha$-coverage is achieved for both groups.
@@ 6.1.1 Network topology and data sources
We simulate a networked learning scenario over a random geometric graph $G=(V,E)$ with $|V|=1024$ nodes and connection radius $r=0.2$. The designated terminal node $t\in V$ collects data from a subset of data sources $B\cup R\subseteq V\setminus\{t\}$. Each of the data sources belongs to one of two demographic groups $B$ or $R$, assigned based on spatial proximity to the terminal (near vs. far).
@@
The algorithms activate edges $A\subseteq E$ to connect data sources to the terminal. For the baselines, the activation cost is defined as the sum of hop-distances from all covered data sources to the terminal. The collected data is used to train a shared MLP classifier. We use a fully connected MLP using an input layer of size $d=2$, followed by two hidden layers with 32 neurons each and ReLU activations, and a final output layer with a single neuron and sigmoid activation for binary classification. We train the model using the Adam optimizer [23] with learning rate $0.01$, batch size $64$, and weight decay $10^{-2}$ for $\ell_{2}$ regularization. The loss is binary cross-entropy. Training runs for 100 epochs with early stopping based on validation loss. The model is trained using the combined dataset collected from data sources temporally connected to the terminal. Accuracy is reported separately on held-out test sets for each group. We also report the coverage of nodes in $B$ and $R$ that can temporally reach $t$. We report the mean values and standard deviations over ten independent runs.
@@ 6.3 Practical scalability
We sampled a connected 20,000-node induced subgraph (201,900 edges) from the Pokec dataset [^1] https://snap.stanford.edu/data/soc-Pokec.html [39]. We randomly selected one node as the terminal and colored 10% of the remaining nodes (balanced 50/50) using the binary gender attribute (Blue = male, Red = female). We then ran FMLbiApprox with full coverage requirement ($\alpha=1.0$) and two precision settings ($\varepsilon=0.01$ and $\varepsilon=0.001$) in 10 independent runs. Table˜2 shows the results. Accuracy represents the ratio of achieved coverage to the required coverage. This real-world experiment confirms the practical scalability and effectiveness of our FMLbiApprox algorithm, which achieves high accuracy under both precision settings.
@@ 7 Limitations
Technical limitations. Our approximation algorithms currently focus on the case of a single terminal and two groups. While this setting already models common use cases, generalizing to arbitrary numbers of terminals and group types remains open. The exact dynamic programming algorithm has a running time complexity in $\mathcal{O}(n^{5})$, limiting scalability. The bicriteria approximation improves efficiency, reducing runtime to $\mathcal{O}(n^{2}+n\varepsilon^{-4}\log^{4}n)$, but trades off exact fairness for scalability: its guarantees degrade with the tree height, leading to a $(1+\varepsilon)^{\log n}$ factor in group coverage.
@@ 15     Keep only the label with minimum $c$ for each $(b,r)$ in $\mathrm{Labels}(v)$
Appendix A Omitted Proofs
@@
The optimal weighted cost on the tree, $C_{T,\text{opt}}$, must satisfy $C_{T,\text{opt}}\leq C_{T}(\lambda_{G}^{*})$, as $\lambda_{G}^{*}$ provides one way to achieve the required reachability on $T$.
Our tree algorithm finds a solution $\lambda_{T}$ with weighted cost $C_{T}$ such that $C_{T}\leq C_{T,\text{opt}}$, either exactly, or via the approximation. Therefore, it follows $|\lambda_{G}|\leq C_{T}\leq C_{T,\text{opt}}\leq C_{T}(\lambda_{G}^{*})$.